A girl is now one-fourth as old as her father, and in seven years, she will be one-half as old as her father was twelve years ago. What are her and her father's present ages?
Answer by richwmiller(17219)

girl=g
father=f
g=f/4
g+7=(f-12)/2
f/4+7=(f-12)/2
multiply both sides by 4
f+28=2f-24
subtract f from both sides
28=f-24
52=f
13=g
check
13=52/4
13=13
13+7=(52-12)/2
20=40/2
20=20